A car traveled 175 miles at a certain speed. If the speed had been 20 mph faster, the trip could of been made in one less hour. Find the speed of the car.
I have both the distances as 175 i have one rate as r and the other rate as r+20 then the car with the rate of r and have as t time and the car with the rate of r+20 i have the time a t-1 is this right?

Answer by stanbon(75887)

A car traveled 175 miles at a certain speed. If the speed had been 20 mph faster, the trip could of been made in one less hour. Find the speed of the car.
---
Slower speed DATA:
distance = 175 miles ; rate = r mph ; time = 175/r hrs.
----------------------------------------------
Faster speed DATA:
distance = 175 miles ; rate = r+20 mph ; time = 175/(r+20) hrs
----------------------------------------------
Equation:
slow time - fast time = 1 hr
175/r - 175/(r+20) = 1
---
Multiply thru by r(r+20):
175(r+20) - 175r = r(r+20)
175*20 = r^2 + 20r
--------------------------
r^2 + 20r - 175*20 = 0
---
(r+70)(r-50) = 0
Positive solution
rate = 50 mph (slower rate)
r+20 = 70 mph (faster rate)
================================
Cheers,
Stan H.
================================




I have both the distances as 175 i have one rate as r and the other rate as r+20 then the car with the rate of r and have as t time and the car with the rate of r+20 i have the time a t-1 is this right?